OpenStudy (cybershadow):

Any1 Help!? Ok so i was working on question

OpenStudy (cybershadow):

In which following function was given \[f(x) = a_{n} + \sin(\pi)X for x \in [2n ,2n+1]\]

OpenStudy (cybershadow):

and f(x) = \[b _{n} + \cos (\pi) x for x \in (2n-1,2n)\]

OpenStudy (cybershadow):

now while going through the solution in the book , i came across this \[f(2n+1)^{+}) = b _{n+1}-1 \]

OpenStudy (cybershadow):

i dint get it , How is RHL of f(x) approaches 2n+1 is \[b _{n+1}-1\]

OpenStudy (cybershadow):

i dint get the \[b _{n+1}\] Part , How come we get this?

OpenStudy (cybershadow):

@Loser66

OpenStudy (cybershadow):

This is the ^ link to the whole question

OpenStudy (cybershadow):

@jdoe0001

OpenStudy (cybershadow):

@radar

OpenStudy (cybershadow):

@skullpatrol

OpenStudy (agent47):

If you're wondering where that +1 comes from, it's because Cos(2kpi) where k is an integer is always 1.

OpenStudy (cybershadow):

actually i got that i m wondering abt bn+1 part

OpenStudy (agent47):

Oh you mean \[b_{n+1}\]

OpenStudy (cybershadow):

Yes!

OpenStudy (agent47):

Ah I see.. They're checking the last endpoint, x = 2n+1

OpenStudy (agent47):

the first part is fine:\[a_n+Sin\pi x\]is defined on [2n, 2n+1]

OpenStudy (cybershadow):

yes, I m confused abt the second part :(

OpenStudy (agent47):

\[b_n+Cos\pi x\]is only defined on (2n-1, 2n), I think they're just incrementing everything for that reason

OpenStudy (agent47):

I mean the ns

OpenStudy (cybershadow):

hmm, yea but why increment n by 1. as it is defiend on (2n-1,2n) so if we have to increment by 1 that wud be for 2n , for 2n+1 we should be incrementing by 2? idk i m confused

OpenStudy (agent47):

something like:\[b_{n+1}+Cos\pi x\]being defined on: (2n-1+1, 2n+1)

OpenStudy (cybershadow):

hmmm

OpenStudy (cybershadow):

sounds legit

OpenStudy (agent47):

or at least that's what i think they're doing. they don't need to add 2, as that's not required by the question

OpenStudy (cybershadow):

but RHL of 2n+1 wud be value greater than 2n+1 , it doesnt contain in (2n-1+1,2n+1) interval, what do u think?

OpenStudy (cybershadow):

for LHL we can use (2n-1+1,2n+1) , but for RHL that interval would not be valid, dont u think so?

OpenStudy (agent47):

true...

OpenStudy (linda3):

i think u have a mod helping you- your fine :P

OpenStudy (phi):

 First, these functions abut each other as we increase n, alternating a+ sin, b+cos,